OpenStudy (anonymous):

A cylinder is full at 385 cm3 and has a radius of 4.5 cm. It currently contains 225 cm3 of water What will the difference in the height of the water be when the cylinder is full? Use 3.14 to approximate pi, and round your answer to one decimal place. cm @triciaal @andrewhaze @linn99123

OpenStudy (anonymous):

@princeharryyy

OpenStudy (princeharryyy):

do u kno the formula for volume of a cylinder

OpenStudy (princeharryyy):

pie*r^2 *h

OpenStudy (princeharryyy):

pie * 4.5^2 * h1 = 385, pie * 4.5^2 * h2 = 225 find h1 and h2 from these equations and finally , h1 -h2 will be your answer

OpenStudy (princeharryyy):

h1 =6.054, h2 = 3.54, (h1-h2 is ur answer)

OpenStudy (princeharryyy):

2.514